Please helpe me solve this math problem:
Cheryl is thirteen years more than twice as old as Brad. In 11 years the sum of thire ages will be 92. How old is Cheryl now?

Found 2 solutions by Earlsdon, oberobic:
Answer by Earlsdon(6294)

Let C = Cheryl's present age and B = Brad's present age.
From the problem description, you can write:
1) C = 2B+13 "Cheryl is 13 years more than twice as old as Brad"
2) (C+11) + (B+11) = 92 "In 11 years the sum of their ages will be 92"
Substitute the C= 2B+13 from equation 1) into equation 2) and solve for B.
2a) (2B+13)+11 +B+11 = 92 Simplify.
3B+35 = 92 Subtract 35 from both sides.
3B = 57 Divide both sides by 3.
B = 19 This is Brad's present age. Use equation 1) to find Cheryl's present age.
C = 2B+13
C = 2(19)+13
C = 38+13
C = 51 This is Cheryl's present age.

Answer by oberobic(2304)

Brad is the youngest, so let his age be x. That means Cheryl is 2x+13 years old.
In 11 years Brad will be x + 11 years old. Cheryl will be 2x+13+11=2x+24 years old. We also know their ages will sum to 92 at that time, so:
x+11+2x+24=92
Simplifying we have:
3x+35=92
Subtracting 35 from both sides:
3x=57
Dividing by 3:
x=19
So, Brad is 19 and Cheryl is 2(19)+13=38+13=51.
In 11 years, Brad will be 30 and Cheryl will be 62.
Checking, we findthat 30+62=92, which is the required check.